# NoTeeline: Supporting Real-Time Notetaking from Keypoints with Large Language Models

### v1

@@ Abstract.
Video has become a popular media form for information sharing and consumption. However, taking notes while watching a video requires significant time and effort. To address this, we propose a novel interactive system, NoTeeline, for taking real-time, personalized notes. NoTeeline lets users quickly jot down key points (micronotes), which are automatically expanded into full-fledged notes that capture the content of the user’s micronotes and are consistent with the user’s writing style. In a within-subjects study (N=12), we found that NoTeeline helps users create high-quality notes that capture the essence of their micronotes with a higher factual correctness (93.2%) while accurately reflecting their writing style. While using NoTeeline, participants experienced significantly reduced mental effort, captured satisfactory notes while writing 47% less text, and completed notetaking with 43.9% less time compared to a manual notetaking baseline.
@@
Figure 1. NoTeeline allows users to jot down key points (micronotes), which are automatically expanded into full notes consistent with the user’s writing style. The input and output are marked in purple and green, respectively. A user writes a (a) micronote while watching a video. NoTeeline captures the timestamp, retrieves the corresponding (b) video transcript, and expands the micronote into (1) full notes. During the expansion, NoTeeline uses the user’s (c) previous notes to infer their writing style. NoTeeline also generates (2) personalized cue questions, (3) summary, and (4) themes for note categorization.
@@ 1. Introduction
Online video platform has emerged as a popular form of communication and information sharing (Tversky et al., 2002; Bétrancourt and Benetos, 2018). While accessing video content is easier than ever, the difficulty lies in capturing notes and organizing key information from these videos (Tabbers et al., 2004; Boyle, 2007). Effective note-taking is crucial for retaining and processing information, yet many users struggle with this task due to dual demands of absorbing new content and creating comprehensive notes simultaneously (Piolat et al., 2005; Jansen et al., 2017; Piolat, 2007). This often results in incomplete or poor-quality notes, or causes users to abandon note-taking altogether (Kafipour and Jafari, 2021; Flanigan and Titsworth, 2020; McCoy, 2020; Bui et al., 2013). By empowering users to take notes more easily, we can help them better engage with the video content.
Current note-taking tools often fall short in addressing the ongoing challenges users face, as they either require manual effort or generate generic, automated notes. Active note-taking tools have explored various methods to assist users in writing and organizing manual notes more efficiently (Subramonyam et al., 2020; Chiu et al., 1999; Cao et al., 2022). Despite offering features that speed up the process, these tools still necessitate manual input for creating comprehensive notes, which remains a significant challenge (Piolat et al., 2005). On the other hand, recent advancements in language processing and Large Language Models (LLMs) have led to tools that automatically generate notes or summaries from videos and live lectures (NoteGPT.io, 2024; Notion, 2024; Otter.ai, 2023). However, these automated tools often lack personalization and do not cater to individual user preferences (Biermann et al., 2022). This creates a gap between the manual, labor-intensive process of traditional note-taking and the impersonal nature of automatic note generation.
In this work, we aim to bridge these gaps by designing NoTeeline, a system that applies the principle of micronotes (Lin et al., 2004) to optimize the note-taking process. NoTeeline enables users to take less demanding micronotes with minimal effort and automatically expand them into full-fledged notes. By Micronotes, we refer to scribbles composed of couple-of-words and informal notations people typically use in quickly writing things down (Lin et al., 2004; Majewski, 2023). These micronotes are concise, causing minimal viewing distraction while capturing essential information densely. We developed a proof-of-concept interface, NoTeeline (Figure 1) that introduces a novel pipeline to enable users to write such micronotes and accurately capture the relevant context in the video to create the full (expanded) note. Micronotes being extremely brief, we need additional context to expand them properly (Monty, 1986). NoTeeline efficiently achieves this by utilizing the video transcript and previous examples of user writeup to decode the micronotes properly. Through a user study, we find that NoTeeline enables significantly faster and easier composition, and synthesis of notes than traditional notetaking tools. This work makes the following contributions:
@@
    We designed NoTeeline, a notetaking tool built around the concept of micronotes. NoTeeline enables users to write quick notes and then utilizes LLMs with relevant context to expand micronotes into full-fledged notes.
@@
    In-depth qualitative and quantitative evaluations that show the strengths and limitations of NoTeeline. We found that NoTeeline generates accurate notes with 93.2% factual correctness while significantly reducing the writing time up to 43.9%.
  3. (3)
    Based on our evaluation and user feedback, we provide design implication for designing better tools for effective Human-AI collaboration .
@@ 2.1. Tools for Supporting Active Notetaking
Active notetaking has been long practiced as an efficient strategy to store and retain information (Peper and Mayer, 1978; Bauer and Koedinger, 2007). Past research has introduced several interactive notetaking tools to support live-notes writing (Weher and Poon, 1994; Hinckley et al., 2012, 2007; Liu et al., 2019; Tashman and Edwards, 2011; Tsai et al., 2024), inking (Forbus et al., 2011; Zheng et al., 2021; Cao et al., 2022), information scaffolding (Bernstein et al., 2008; Swearngin et al., 2021) and note organization (Spoerri, 2002; Meng et al., 2016; Subramonyam et al., 2020; He, 2014). NoteLook supports multimedia note taking in meetings in an integrated conference room (Chiu et al., 1999). VideoSticker proposes a visual notetaking tool for lecture videos utilizing object symbols to create multimodal, non-linear notes (Cao et al., 2022). AQuA introduces a question-answering interface that helps users to watch software tutorial videos to get quick information (Yang et al., 2024). Livenote allows students to take lecture notes and annotate lecture slides cooperatively (Kam et al., 2005). NoteLink allows students to link their hand-written notes with the relevant video so that they can replay it easily (Srinivasa et al., 2021). Although these tools help users to organize and link source materials faster, users still have to write the whole note manually, often requiring additional laborious revisions and manual effort, causing significant cognitive barrier (Piolat et al., 2005). On the contrary, NoTeeline focuses on automatically generating the text of full notes from micronotes to shift effort from text notation to capturing key ideas during active notetaking with users’ own note points. Instead of focusing on copying information word-for-word, users can concentrate on understanding and capturing the core ideas, minimizing the cognitive load associated with manual note-taking.
@@ 2.2. Supporting Shorthand Writing
Shorthand writing techniques have been popular with journalists for centuries (McCay, 2021), to support higher writing speeds without loss of information. Inspired by their success, many digital inking software drew inspiration from shorthand techniques (Leedham et al., 2009). This paved the way for a large body of research for gesture keyboard (Zhai and Kristensson, 2012). SHARK augments stylus keyboarding with shorthand gesturing by assigning a shorthand symbol to each individual word (Zhai and Kristensson, 2003; Kristensson and Zhai, 2004). Swipeboard encodes a sequence of gesture input into unique representations of each character to support faster typing in ultra-small devices such as smart watches (Chen et al., 2014). Both Swipeboard and SHARK report a higher writing speed compared to traditional keyboard. Padilla et al. introduce a deep learning-based methodology to automatically convert Gregg shorthand to English words (Padilla et al., 2020). Although these tools show the potential of shorthand-based writing techniques in the modern era, they specifically focus on gesturing interaction. Oftentimes, they also require users to learn existing shorthand symbols. NoTeeline provides users the flexibility to use their own abbreviated words and symbols without requiring users to learn new gestures to use our system. To be best of our knowledge, NoTeeline is the first tool to support micro-notetaking from videos and easy expansion of micronotes into full notes that are consistent with users’ corresponding writing styles.
@@ 2.3. Supporting Writing with LLMs
LLMs have emerged as popular tools for writing assistance because of their versatile generation capabilities. A wide body of research has focused on enhancement of LLMs as writing assistant through building novel interaction tools (Chen et al., 2019; Clark et al., 2018; Freiknecht and Effelsberg, 2020; Osone et al., 2021; Chung et al., 2022; Ghai and Mueller, 2021) and improving generation capacity of these models (Mysore et al., 2023; Li et al., 2023; Trajanovski et al., 2021; Tang et al., 2017a). Wordcraft allows users to write stories with LLM collaboratively with advanced features such as replacement or further prompting (Yuan et al., 2022). Ramblers allows users to create voice memos and collaboratively use LLM to edit and expand on the voice memos to create a gist (Lin et al., 2024). Pearl proposes a retrieval-augmented pipeline that utilizes previous user-authored documents to generate texts of specific user style (Mysore et al., 2023). ExpaNet retrieves relevant documents to an input query and expands it using the retrieved documents (Tang et al., 2017b). Outside academia, there is a growing number of commercially available tools for personal writing assistant as well (ProductHunt, 2024). While people frequently use these tools for assisting writing (Lee et al., 2022; Yuan et al., 2022), LLMs may struggle to preserve the style of experienced writers (Ippolito et al., 2022). Current research is increasingly directed towards personalized writing aids (Biermann et al., 2022). Compared to the above-mentioned tools, NoTeeline is designed for robust user customization of note generation from start to end via control of references (i.e: micronote) and writing style.
@@ 3. User Study Setup
Our study consists of two phases. First, we performed a needfinding study with 12 participants (§4) to understand the struggles they face in writing full notes in existing notetaking tools. Next, participants evaluated NoTeeline in one hour session. (§7). The entire session was recorded in Zoom. We recruited our participants through mailing lists and announcement boards. The participants are situated in North America. They received a \$30 Amazon gift card for their participation. The study was approved by the IRB committee in our home institution.
@@ 4.1. Procedure
We conducted a needfinding study to understand the current challenges users encounter with their current note-taking process. Specifically, we want to uncover underlying factors that make writing the full note difficult. We also want to understand whether users currently use AI-assistance for notetaking and explore the factors impacting their trust in AI generated notes. The results of this study informed the design choices we made for NoTeeline. Each session were approximately 20-30 minutes long.
@@ 4.2. Insights and Observations
Our participants reported using both manual note taking tools as well as AI assisted tools. The most common app used by participants was the built-in Macbook Notes app (n=9). Many participants use pen-and-paper notes and later digitize them manually. In terms of AI-based tools, some participants used ChatGPT (n=4) for summarizing meetings and video lectures. Others used Notion AI, Figjam AI, and Gemini plugin in Google Docs for specific usecases. Participants highlighted several challenges they face with their current note taking process. We summarize these findings into three key themes through iterative rounds of coding guided by qualitative data analysis (Yin, 2015):
@@
C1. Users struggle to notate all desired information while taking notes from audio, video, and other real-time sources (×7)($\times$ 7)( × 7 ). Users mention resource constraints–time, attention, and mental load–in the practice of notetaking. P8 explains how their handwriting speed cannot match a professor’s lecturing pace. P7 details the additional difficulties imposed by being a non-native English speaker.
@@
P11 finds it challenging to take notes while listening to podcasts, especially during physical activities like running on a treadmill. They describe the cumbersome process to jot down : ‘I have to slow down.. and then type it down in [a few words]… then I again need to speed up the treadmill and continue.’ The user emphasizes how the ad hoc nature of this process complicates effective note-taking, making it difficult to capture details in real-time.
@@
C2. Lack of agency in AI-assisted notetaking tools (×$\times 6$). Users expressed the desire to maintain agency over AI assisted notetaking tools. This is important given concerns about LLM ‘hallucinating’ non-existent content (Ji et al., 2023). P1 says ‘I would, at the end of the day, want some kind of source of truth… based on what the notes are going to be’. They reported to prefer using automated note features ‘as a compliment to whatever that [they’re] using’ while ‘always going to take [their] notes no matter how’, because of the lack of trust in it. P2 further explains her thought - ‘I hope it’s a tool for me… to help me get better understanding of the video, not a tool for me to tell me what I should know, because I still want myself to have the control of what the tool is doing.’
@@
C3. Complex features and workflow of current tools. (×7)($\times$ 7)( × 7 ). Participants highlighted that the complex features and fragmented workflows of current note-taking tools often hinder usability. One common challenge involved managing two separate applications—one for watching videos and another for taking notes—which made the process cumbersome. Many expressed a preference for integrating note-taking features directly within the video viewport, facilitating smoother interactions and easier linking of notes to specific video segments. This integration would also simplify the task of reviewing notes later on.
@@
Furthermore, participants reported that the overwhelming nature of feature-rich tools like Notion discouraged full utilization. As P5 described, ‘I don’t end up using the extent of features that Notion has to offer… it’s so overwhelming .. so I think [I prefer] a simple interface but with the ability that as you go forward, you can start to explore more’. In other words, the features are not well-situated or contextualized within the workflow, and presenting them all at once can overwhelm users, preventing them from fully utilizing the benefits. In fact, it can deter users from engaging with the features altogether.
5. NoTeeline Interface
Figure 2. The NoTeeline user interface consists of (a) the Cues panel (outlined in blue) for review questions generated based on the notes, (b) the Notes panel (outlined in orange) for user notetaking and note expansion, and (c) the Summary panel (outlined in purple) for note summary. The layout is inspired by the Cornell method, as seen in the side-by-side comparison.

### v2

# NoTeeline: Supporting Real-Time, Personalized Notetaking with LLM-Enhanced Micronotes

@@ Abstract.
Taking notes quickly while effectively capturing key information can be challenging, especially when watching videos that present simultaneous visual and auditory streams. Manually taken notes often miss crucial details due to the fast-paced nature of the content, while automatically generated notes fail to incorporate user preferences and discourage active engagement with the content. To address this, we propose an interactive system, NoTeeline, for supporting real-time, personalized notetaking. Given ‘micronotes’, NoTeeline automatically expands them into full-fledged notes using Large Language Model (LLM). The generated notes build on the content of micronotes by adding relevant details while maintaining consistency with the user’s writing style. In a within-subjects study (n=12), we found that NoTeeline creates high-quality notes that capture the essence of their micronotes with 93.2% factual correctness and accurately align with their writing style (8.33% improvement). Using NoTeeline, participants could capture their desired notes with significantly reduced mental effort, writing 47.0% less text and completing their note in 43.9% less time compared to a manual notetaking baseline. Our results suggest that NoTeeline enables users to integrate LLM assistance in a familiar notetaking workflow while ensuring consistency with their preference.
@@
Figure 1. NoTeeline enables users to take notes while watching videos on-the-fly by automatically converting brief micronotes into fully-expanded notes in their writing style. While watching a video, the user jots down quick keypoints (a). NoTeeline retrieves the transcript at the timestamp (b). In addition, NoTeeline references a small collection of the user’s previous notes (c). Given (b) and (c), NoTeeline automatically expands the micronote into a complete note in their writing style (1). Inspired by the Cornell notetaking system (Saran et al., 2022), NoTeeline generates review questions (2), a brief summary (3), and groups notes into key themes (4) to assist users in note review.
@@ 1. Introduction
Effective notetaking is crucial for capturing and organizing information, allowing users to record essential details for later review. Whether in academic, professional, or personal settings, notes help individuals track important points, revisit information when needed, and maintain focus on the task at hand without getting overwhelmed by the flow of information. However, notetaking can be very challenging in a fast-paced environment, such as when watching information-dense videos or videos that are difficult to pause. Users may struggle with the high mental demands, forcing them to divide their attention between absorbing new content and documenting it (Piolat et al., 2005; Jansen et al., 2017; Piolat, 2007). For example, when taking notes from information-dense videos, people often pause the video frequently in order to write notes, which disrupts the natural flow of content engagement (Cao et al., 2022) and forces the user to dedicate mental and physical effort to the mechanical act of notetaking. The inability to pause (e.g. in some live-streamed content or content on platforms that lack full media controls) adds even more pressure, often resulting in incomplete notes that may be difficult to interpret later (McCoy, 2020; Kafipour and Jafari, 2021; Flanigan and Titsworth, 2020)
@@
With the emergence of Large Language Models (LLMs,) people have explored building automated notetaking tools, such as NoteGPT (NoteGPT.io, 2024; Notion, 2024; Otter.ai, 2023). These tools typically work by automatically generating notes or summaries from the video transcript or live lecture recordings in one go. Thus, the generated notes lack personalization (Biermann et al., 2022). For example, the notes may be too general and not tailored to the user’s specific needs. A user may only be interested in specific topics of the material, or they might prefer their notes to be written in a particular style that reflects how they process information. Automated notes can feel impersonal and robotic (Biermann et al., 2022).
Prior studies have observed that, in fast-paced notetaking situations, people often resort to shorthand techniques like micronotes(Lin et al., 2004). For example, they may jot down short keypoints containing shortened words or informational notations in order to write things down more quickly (see Figure 1a). However, while micronotes help users keep up with the pace, they are often too brief or vague to be useful during the later review. As a result, these micronotes are often abandoned and not revisited (Kleek et al., 2009). Thus, we observed a gap between quickly capturing information and having detailed, meaningful notes that can be revisited with clarity.
In this work, we leverage the practice of micro-notetaking and build an assistive tool around this practice, specifically for videos. Video content presents a unique challenge for notetaking due to its temporal nature and often being information-dense, which leaves little time for users to capture important details. We introduce NoTeeline  [^1] NoTeeline’s name is inspired by Teeline(Contributors, 2024), a popular shorthand technique., an interactive notetaking assistance tool that aims to automatically expand user micronotes into complete notes in real-time that also aligns with their personal writing style. We propose a novel pipeline that leverages a window of the video transcript as well as a handful of examples of the user’s previous notes. Through a within-subjects user study, we found that NoTeeline helps users effectively capture notes with minimal disruption (e.g., reduced video pauses), capturing more content using fewer words and shorter time compared to a manual notetaking baseline. Participants rate the generated note expansions as high quality (average 6.0 on 7-point Likert scale, higher is better), align with the participants’ micronote content (relatively high semantic similarity between micronotes and expanded notes measured with SBERT score (Reimers and Gurevych, 2019)), are consistent with participants’ writing style (8.33% improvement in chi-squared Distance (Kilgarriff, 2001)), and the notes contain low hallucination (93.2% factual correctness score measured with HHEM score (Vectara, 2024)).
In summary, this work makes the following contributions.
@@
    We designed NoTeeline, an interactive notetaking tool built around the concept of micronotes. NoTeeline helps users write quick notes in fast-paced contexts and utilizes LLMs to expand micronotes into full-fledged notes.
@@
    We conducted a within-subjects user study to evaluate the performance and utility of NoTeeline. Our quantitative and qualitative results show that with NoTeeline, users can take high-quality notes with higher efficiency and satisfaction, compared to a manual notetaking baseline.
@@ 2.1. Tools for Supporting Active Notetaking
Active notetaking has been long practiced as an efficient strategy to store and retain information (Peper and Mayer, 1978; Bauer and Koedinger, 2007). Past research has introduced several interactive notetaking tools to support live-notes writing (Weher and Poon, 1994; Hinckley et al., 2012, 2007; Liu et al., 2019; Tashman and Edwards, 2011; Tsai et al., 2024), inking (Forbus et al., 2011; Zheng et al., 2021; Cao et al., 2022), information scaffolding (Bernstein et al., 2008; Swearngin et al., 2021) and note organization (Spoerri, 2002; Meng et al., 2016; Subramonyam et al., 2020; He, 2014). NoteLook supports multimedia note taking in meetings in an integrated conference room (Chiu et al., 1999). VideoSticker proposes a visual notetaking tool for lecture videos utilizing object symbols to create multimodal, non-linear notes (Cao et al., 2022). AQuA introduces a question-answering interface that helps users to watch software tutorial videos to get quick information (Yang et al., 2024). Livenote allows students to take lecture notes and annotate lecture slides cooperatively (Kam et al., 2005). NoteLink allows students to link their hand-written notes with the relevant video so that they can replay it easily (Srinivasa et al., 2021). Although these tools help users to organize and link source materials faster, users still have to write the whole note manually, often requiring additional laborious revisions and manual effort, causing a significant cognitive barrier (Piolat et al., 2005). On the contrary, NoTeeline focuses on automatically generating the text of full notes from micronotes to shift effort from text notation to capturing key ideas during active notetaking with users’ own note points.
@@ 2.2. Supporting Shorthand Writing
Shorthand writing techniques have been popular with journalists for centuries (McCay, 2021), to support higher writing speeds without loss of information. Inspired by their success, many digital inking software drew inspiration from shorthand techniques (Leedham et al., 2009). This paved the way for a large body of research for gesture keyboard (Zhai and Kristensson, 2012). SHARK augments stylus keyboarding with shorthand gesturing by assigning a shorthand symbol to each word (Zhai and Kristensson, 2003; Kristensson and Zhai, 2004). Swipeboard encodes a sequence of gesture input into unique representations of each character to support faster typing in ultra-small devices such as smart watches (Chen et al., 2014). Both Swipeboard and SHARK report a higher writing speed compared to traditional keyboard. Padilla et al. introduce a deep learning-based methodology to automatically convert Gregg shorthand to English words (Padilla et al., 2020). Although these tools show the potential of shorthand-based writing techniques in the modern era, they specifically focus on gesturing interaction. Oftentimes, they also require users to learn existing shorthand symbols. NoTeeline provides users the flexibility to use their own abbreviated words and symbols without requiring users to learn new gestures to use our system. To be best of our knowledge, NoTeeline is the first tool to support micro-notetaking from videos and easy expansion of micronotes into full notes that are consistent with users’ corresponding writing styles.
@@ 2.3. Supporting Writing with LLMs
LLMs have emerged as popular tools for writing assistance because of their versatile generation capabilities. A wide body of research has focused on enhancement of LLMs as writing assistant through building novel interaction tools (Chen et al., 2019; Clark et al., 2018; Freiknecht and Effelsberg, 2020; Osone et al., 2021; Chung et al., 2022; Ghai and Mueller, 2021) and improving generation capacity of these models (Mysore et al., 2023; Li et al., 2023; Trajanovski et al., 2021; Tang et al., 2017a). Wordcraft allows users to write stories with LLM collaboratively with advanced features such as replacement or further prompting (Yuan et al., 2022). Ramblers allows users to create voice memos and collaboratively use LLM to edit and expand on the voice memos to create a gist (Lin et al., 2024). Pearl proposes a retrieval-augmented pipeline that utilizes previous user-authored documents to generate texts of specific user style (Mysore et al., 2023). ExpaNet retrieves relevant documents to an input query and expands it using the retrieved documents (Tang et al., 2017b). Outside academia, there is a growing number of commercially available tools for personal writing assistants as well (ProductHunt, 2024). While people frequently use these tools for assisting writing (Lee et al., 2022; Yuan et al., 2022), LLMs may struggle to preserve the style of experienced writers (Ippolito et al., 2022). Current research is increasingly directed towards personalized writing aids (Biermann et al., 2022). Compared to the above-mentioned tools, NoTeeline is designed for robust user customization of note generation from start to end via control of references (i.e.: micronote) and writing style.
@@ 3. User Study Setup
Our study consists of two phases. First, we performed a needfinding study with 12 participants (§4, referred to as P1-P12) to understand the struggles they face in writing full notes in existing notetaking tools. Next, participants evaluated NoTeeline in an hour session. (§7). The session was recorded in Zoom. We recruited our participants through mailing lists and announcement boards. The participants are situated in North America. They received a \$30 Amazon gift card for their participation. The study was approved by the IRB committee in our home institution.
@@ 4.1. Procedure
We conducted a needfinding study to understand the current challenges users encounter with their current notetaking process. Specifically, we want to uncover underlying factors that make writing the full note difficult. The results of this study informed the design choices we made for NoTeeline. Each session was approximately 20-30 minutes long.
Our participants reported using both manual notetaking tools as well as AI-assisted tools. The most common app used by participants was the built-in Macbook Notes app ($n=9$). Many participants use pen-and-paper notes ($n=7$) and later digitize them manually. In terms of AI-based tools, some participants used ChatGPT ($n=4$) for summarizing meetings and video lectures. Others used Notion AI, Figjam AI, and Gemini plugin in Google Docs for specific use cases.
@@ 4.2. Insights and Observations
Participants highlighted several challenges they face with their current notetaking process. We summarize these findings into three key themes through iterative rounds of coding guided by qualitative data analysis (Yin, 2015):
@@
C1. Users struggle to capture all desired information while taking notes from audio, video, and other real-time sources $(n=7)$. Users mention resource constraints–time, attention, and mental load–in the practice of notetaking. P8 explains how their handwriting speed cannot match a professor’s lecturing pace. P7 details the additional difficulties imposed by being a non-native English speaker.
@@
P11 finds it challenging to take notes while listening to podcasts, especially during physical activities like running on a treadmill. They describe the cumbersome process to jot down: ‘I have to slow down.. and then type it down in [a few words]… then I again need to speed up the treadmill and continue.’ The user emphasizes how the ad hoc nature of this process complicates effective notetaking, making it difficult to capture details in real time.
@@
C2. Lack of agency in AI-assisted notetaking tools ($n=6$). Users expressed the desire to maintain agency over AI-assisted notetaking tools. This is important given concerns about LLM ‘hallucinating’ non-existent content (Ji et al., 2023). P1 says ‘I would, at the end of the day, want some kind of source of truth… based on what the notes are going to be’. They reported to prefer using automated note features ‘as a compliment to whatever that [they’re] using’ while ‘always going to take [their] notes no matter how’, because of the lack of trust in it. P2 further explains her thought - ‘I hope it’s a tool for me… to help me get a better understanding of the video, not a tool for me to tell me what I should know because I still want myself to have the control of what the tool is doing.’
@@
C3. Complex features and workflow of current tools. $(n=7)$. Participants highlighted that the complex features and fragmented workflow of current notetaking tools often hinder usability. One common challenge involved managing two separate applications—one for watching videos and another for taking notes—which made the process cumbersome. They expressed a desire for having notetaking features within the video application, as this would facilitate a more seamless interaction, especially if the notes are linked to specific video segments to aid in later review.
@@
Furthermore, participants reported that the overwhelming nature of feature-rich tools like Notion discouraged full utilization. As P5 described, ‘I don’t end up using the extent of features that Notion has to offer… it’s so overwhelming .. so [I prefer] a simple interface but with the ability that as you go forward, you can start to explore more’. In other words, the features are not well-situated or contextualized within the workflow, and presenting them all at once can overwhelm users, preventing them from fully utilizing the benefits.